A number minus 12 times its reciprocal is equal to 1. Find the number.
Answer by mananth(16949)

let number be x
x-12*1/x = 1
multiply by x
x^2-12=x
x^2-x-12=0
x^2-4x+3x-12=0
x(x-4)+3(x-4)=0
(x-4)(x+3)=0
x=4 OR -3